determine whether the given ordered pair is a solution to the system.
(-2,-3)
3x + y = -9
2x + 3y = 5

Found 3 solutions by rfer, checkley77, stanbon:
Answer by rfer(10432)

number one=yes, two no

Answer by checkley77(12569)

(-2,-3)
3x + y = -9
3*-2-3=-9
-6-3=-9 OK.
-----------------
2x + 3y = 5
2*-3+3*-3=5
-6-9=5
-15 does not = 5 NOT OK.

Answer by stanbon(48568)

determine whether the given ordered pair is a solution to the system.
(-2,-3)
3x + y = -9
2x + 3y = 5
---------------------
Substitute the x/y values into the equation to see
if you get true statements:
3*-2+-3 = 5
-6-3 = 5
true
-------
2*-2+3*-3 = 5
-4-9 = 5
wrong
-----------------
The pair (-2,-3) is not a solution of the system of equations.
-----------------
Cheers,
Stan H.